OpenStudy (anonymous):

could someone help me Solve 2x2 − 8x = −7.

OpenStudy (anonymous):

i thought the -7 was going to be the c

pooja195 (pooja195):

Add the 7 2x^2-8x-7=0 a=2 b=8 c=7 \[\huge~\rm~x=\frac{ -(8) \pm \sqrt{(8)^2-4(2)(7)} }{ 2(2) }\]

OpenStudy (anonymous):

i simplified it what would i do next to simplify it more @pooja195

pooja195 (pooja195):

Show me what u have

OpenStudy (anonymous):

-8 +sqrt 8/4 or -8 - sqrt 8/4

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):



OpenStudy (anonymous):

none of my answers have a 1/2 in them

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

It would help if i knew what they were

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Im not sure... @EclipsedStar @Data_LG2

OpenStudy (anonymous):

@ganeshie8

eclipsedstar (eclipsedstar):

I tried it out, my answer doesn't match up either...how weird

OpenStudy (mathstudent55):

\(2x^2 − 8x = −7\) \(2x^2 - 8x + 7 = 0\) \(a = 2\); \(b = -8\); \(c = 7\) \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) \(x = \dfrac{-(-8) \pm \sqrt{(-8)^2 - 4(2)(7)}}{2(2)}\) \(x = \dfrac{8 \pm \sqrt{64 - 56}}{4}\) \(x = \dfrac{8 \pm \sqrt{8}}{4}\) \(x = \dfrac{8 \pm 2\sqrt{2}}{4}\) \(x = \dfrac{2 \cdot 4 \pm 2\sqrt{2}}{2 \cdot 2}\) \(x = \dfrac{4\pm \sqrt{2}}{ 2}\) \(x = \dfrac{4}{2} \pm \dfrac{\sqrt{2}}{ 2}\) \(x = 2 \pm \dfrac{\sqrt{2}}{ 2}\)

eclipsedstar (eclipsedstar):

oh I see what you did mathstudent... I only stopped at \(x = \dfrac{4\pm \sqrt{2}}{ 2}\) didn't simplify lol

OpenStudy (mathstudent55):

Where you stopped was a correct answer, but this way you can match one of the choices.